### v2

# GTENN: A Spatiotemporal Graph Neural Framework for Community Discovery in Dynamic Social Networks

Abstract
Community discovery is one of the key issues in the study of dynamic social networks. Traditional community discovery algorithms mainly focus on the formation and dissolution of links between nodes, and thus fail to capture richer spatial and temporal patterns underlying network evolution. To address this limitation, we propose GTENN, a spatiotemporal graph neural framework for community discovery in dynamic social networks. GTENN integrates spatial structure and temporal dynamics within a unified embedding architecture. First, Graph Convolutional Networks (GCN) are employed to aggregate latent spatial information and learn expressive node representations at each snapshot. Next, Gated Recurrent Units (GRU) are used to model temporal evolutions of node embeddings, effectively capturing node dynamism and relationship propagation across time. Finally, a Self-Organizing Map (SOM) is applied to the learned spatiotemporal embeddings to cluster nodes and infer their community affiliations. We conduct experiments on four types of dynamic networks, and the results show that GTENN consistently outperforms traditional community discovery algorithms in terms of purity, normalized mutual information, homogeneity, and completeness. These findings demonstrate the superior ability of GTENN to accurately uncover evolving community structures hidden in dynamic social networks.
keywords:
Community Discovery; Dynamic Social Networks; Spatiotemporal Graph Embedding; Recurrent Unit; Self-organizing Map
1 Introduction
Social networks have become an indispensable part of daily lives, with explosive growth in both scale and complexity. Dynamic social networks are composed of a vast number of user nodes and the evolving connections between them. These networks contain rich information and intricate structural patterns that are worth exploring.
Community discovery[1], one of the key tasks in social network analysis, aims to identify community structures in the network that exhibit strong internal connections while being relatively independent of other groups. The applications of community discovery span a wide range of domains, providing valuable insights and effective solutions. In targeted marketing and personalization, communities offer a deeper understanding of shared interests and behavioral patterns[2], enabling the development of more effective recommendation systems and customized advertising strategies. In epidemiology and public health[3, 4], identified communities can represent populations at higher risk or groups vulnerable to the spread of infectious diseases, facilitating the design of targeted intervention and containment measures. Similarly, in security and fraud detection, community discovery plays a crucial role in identifying anomalous groups within financial or communication networks[5, 6, 7], helping analysts detect fraudulent activities, cyber threats, or other malicious entities. Finally, in social science and behavioral analysis[8, 9], understanding community dynamics provides critical insights into cultural trends, patterns of social influence, and the evolution of social norms over time. These diverse applications underscore the importance of studying communities within dynamic networks to address real-world challenges.
In static social networks, numerous community discovery algorithms[10] have achieved certain results. However, the temporal variability of dynamic social networks poses significant challenges for community discovery[11, 12]. Traditional static algorithms[13] face challenges in capturing dynamic changes in network properties across different time points, and fail to accurately reflect the real community structure and its evolution process. Recently, graph embedding techniques[14] emerged in the field of social network analysis. Such techniques map nodes in the network to a low-dimensional vector space, thereby preserving the structure and semantic information of nodes. However, most existing graph embedding methods still focus primarily on static networks, with significant limitations in the joint processing of temporal and spatial dimension information in dynamic social networks[15].
To address these issues, in this paper, we propose a community discovery algorithm based on spatiotemporal graph embedding. The proposed algorithm fully accounts for the spatial connectivity relationships and temporal trends of the nodes in dynamic social networks. Through innovative embedding techniques, it integrates spatiotemporal information into a unified framework, enabling more accurate detection of the community structure of networks and a clearer visualization of their evolution path. This approach provides strong support for a deeper understanding of the inherent laws and mechanisms of dynamic social networks and has great potential for applications in social network analysis, recommendation systems, information dissemination, and other fields.
This paper is organized as follows. In Sect. 1, we briefly discuss the motivation to propose a dynamic community discovery algorithm. In Sect. 2, we review some of the progress and limitations of current community discovery algorithms. In Sect. 3, we discuss the necessary definitions related to GTENN and introduce the algorithm. Experiments and comparisons are conducted in Sect. 4. And finally Sect. 5 concludes the paper.
2 Related Works
Researchers have made significant progress in the study of community discoveries within dynamic social networks. [16] proposed a dynamic community discovery algorithm based on spectral clustering (SC-DCDA) to address the problems of network structure analysis, insufficient clustering information, and excessive computational overhead caused by the lack of historical information necessary for community nodes in static community discovery algorithms for social networks. This algorithm is based on the division of community network evolution events and the analysis of major community events. Traditional community discovery methods assume that nodes have an equal status in the network, ignoring the influence of nodes in community formation. To address this issue, [17] proposed a community discovery method based on node influence expansion. This approach begins by calculating the local influence of the nodes using Monte Carlo simulation. Next, a novel eccentricity-based method is introduced to screen bridge nodes and enhance seed quality. Finally, dynamic programming is applied to expand the seeds and generate communities. To reveal hidden dynamic phenomena in virtual social networks, [18] proposed a multi-time dense subgraph discovery algorithm for online social network discovery, which considers both topic and time. In this algorithm, conversations or comments are modeled as an interactive network, which is then partitioned into topic-specific communities using its topological structure. Within each community, dense subgraphs are mined on the basis of time dimension, capturing temporal dynamics.
However, current dynamic social network community discovery algorithms still have many limitations: (a) The explosive growth of social network data increases computational complexity, memory usage, and runtime. Many algorithms require recalculating at each time step, causing the time complexity to rise significantly with the number of steps, hindering real-time performance. (b) Some algorithms neglect temporal correlations and fail to capture the true evolution of communities and complex structures such as nesting, overlap, and hierarchy, leading to the loss of critical information. (c) Algorithms often struggle with the dynamic nature of social networks, exhibiting poor adaptability to structural changes and sensitivity to parameter adjustments. In addition, they are vulnerable to noise and abnormal data, compromising accuracy measured in purity and NMI and reliability. Considering these limitations, to fully extract the dynamic features of the network and the interaction information between nodes, and to deeply mine the dynamic features of the network, we propose a community discovery algorithm based on spatiotemporal graph embedding (GTENN). This algorithm employs Graph Convolutional Neural Networks (GCNs) to aggregate latent spatial information, effectively representing the embedding of nodes in space. The temporal evolutions of nodes are then modeled using Gated Recurrent Units (GRU), thereby solving problems such as node dynamism and relationship transmission. Finally, a Self-Organizing Map (SOM) is applied to cluster dynamic network representations and identify community affiliations of nodes.
3 Dynamic Community Discovery Algorithm
In this section, we first provide definitions and descriptions of the relevant concepts used in this paper. Then, we introduce the proposed algorithm.
3.1 Definitions
Definition 1 (Dynamic Network).
Let $V=\{v_{1},v_{2},\dots,v_{n}\}$ be the set of nodes and $E=\{e_{ij}\mid i\neq j,i,j\in\{1,\dots,n\}\}$ the set of potential edges in a dynamic network. At each time step $t$, the snapshot network is defined as $g_{t}=(V,E_{t})$, where $E_{t}\subseteq E$ varies over time while the node set $V$ remains fixed.
The full dynamic network is then represented as a sequence of time-varying snapshots:
$G=\{g_{1},g_{2},\dots,g_{T}\},\quad 1\leq t\leq T.$
Each snapshot $g_{t}$ is associated with an adjacency matrix $A_{t}\in\mathbb{R}^{n\times n}$, where $A_{t}(i,j)$ encodes the edge weight from node $v_{i}$ to node $v_{j}$ at time $t$. Unless otherwise specified, all graphs are assumed to be undirected, i.e., $A_{t}=A_{t}^{\top}$, and unweighted, i.e., $A_{t}(i,j)\in\{0,1\}$.
Definition 2 (Dynamic Community).
We use $C_{t}=\{c_{t}^{1},c_{t}^{2},\dots,c_{t}^{p}\mid\forall t\in\{1,\dots,T\}\}$ to denote the $p$ communities at time slice network $g_{t}$, where each $c_{t}^{i}\subseteq V$ is a subset of nodes belonging to the $i$-th community at time $t$.
The dynamic community structure over the entire dynamic network $G$ is represented as:
$C=\{C_{1},C_{2},\dots,C_{T}\}.$
Definition 3 (Graph Embedding).
Graph embedding is a mapping function $\Gamma$ that projects each node $v_{i}$ at time $t$ into a low-dimensional latent space:
$\Gamma:v_{i}^{t}\mapsto f_{i}^{t}\in\mathbb{R}^{d},\quad 1\leq i\leq n,\;1\leq t\leq T,\;d\ll n.$
Here, $f_{i}^{t}$ is the embedding vector for node $v_{i}$ at time $t$, and $d$ is the embedding dimension. The matrix of all node embeddings at time $t$ is denoted by $F_{t}=[f_{1}^{t},f_{2}^{t},\dots,f_{n}^{t}]^{\top}\in\mathbb{R}^{n\times d}$, and the full sequence of embedding matrices is denoted as:
$F_{S}=\{F_{1},F_{2},\dots,F_{T}\}.$
Definition 4 (Adjacency Matrix).
Given a dynamic network $G=\{g_{t}\}_{t=1}^{T}$, each snapshot $g_{t}=(V,E_{t})$ consists of a fixed set of nodes $V$ and a set of edges $E_{t}$ at time $t$. The adjacency matrix at time $t$ is defined as
$A_{t}(i,j)=\begin{cases}1,&\text{if }(v_{i},v_{j})\in E_{t},\\ 0,&\text{otherwise},\end{cases}\quad\forall v_{i},v_{j}\in V.$
Since we consider undirected graphs [^1] Self-loops are not included unless explicitly added for normalization purposes., the adjacency matrix is symmetric, i.e., $A_{t}(i,j)=A_{t}(j,i)$.
Symbol Description
$G$ Dynamic network (sequence of graphs)
$g_{t}$ Graph snapshot at time $t$
$A_{t}$ Adjacency matrix of $g_{t}$
$\hat{A}_{t}$ Normalized adjacency matrix at time $t$
$t$ Current time step
$T$ Total number of time steps
$n$ Number of nodes
$m$ Number of edges
$d$ Dimension of embedding vector
$C$ Dynamic community collection $\{C_{t}\}_{t=1}^{T}$
$F_{t}$ Embedding matrix at time $t\in\mathbb{R}^{n\times d}$
$F_{S}$ Sequence of all embeddings over time
$f_{i}^{t}$ Embedding vector of node $v_{i}$ at time $t$
$\Gamma$ Embedding function
Table 1: Notations used throughout the paper, including dynamic network representations, time-indexed graph structures, embedding dimensions, and community-related variables.
3.2 The GTENN Algorithm
The dynamic network community discovery process based on spatiotemporal graph embedding, shown in Fig. 1, consists of three main components: spatial feature extraction, temporal feature extraction, and community discovery. In order to fully utilize the spatial information of the nodes in the network, we first employ GCNs to learn the feature representation of the nodes from a spatial perspective and aggregate the social information of users who communicate with each other. Next, we share the GCNs’ weight parameters using GRU to maintain the dynamics of spatial nodes and extract the features of dynamic nodes, addressing node dynamics and long-term dependencies to form the feature representation of the dynamic network. Finally, the SOM algorithm is applied to analyze the network embedding and perform community discovery for dynamic artificial and real social networks. We dub this process as the GTENN algorithm.
Figure 1: Framework of the proposed GTENN algorithm. The process integrates three main components: (i) spatial feature extraction using GCNs to learn node representations from network topology, (ii) temporal feature extraction via GRU to capture node dynamics and long-term dependencies by sharing GCN weights across time, and (iii) community discovery using SOM to analyze the learned embeddings and identify evolving communities in dynamic networks.
3.3 Spatial Feature Extraction
In the spatial feature extraction stage, we start with node interaction learning and use GCNs to learn the topological structure information of nodes in the snapshot network and obtain the embeddings of nodes. This process involves two key aspects: space embedding learning and weight parameter sharing.
3.3.1 Space Embedding Learning
GCNs consist of multiple layers of graph convolutions that aggregate neighborhood features. At time step $t$, in the $l$-th GCN layer, the model takes as input the node representation matrix $F_{t}^{l-1}\in\mathbb{R}^{n\times d}$, the snapshot adjacency matrix $A_{t}\in\mathbb{R}^{n\times n}$, and a learnable weight matrix $W_{t}^{l-1}\in\mathbb{R}^{d\times d^{\prime}}$. The updated node representation matrix is computed as:
$F_{t}^{l}=\Gamma\left(F_{t}^{l-1},A_{t},W_{t}^{l-1}\right)=\sigma\left(\hat{A}_{t}F_{t}^{l-1}W_{t}^{l-1}\right),$ (1)
where $\sigma(\cdot)$ is the activation function, and in this paper we adopt the LeakyReLU function. The matrix $\hat{A}_{t}\in\mathbb{R}^{n\times n}$ is the symmetrically normalized adjacency matrix used as a graph convolution filter, computed as follows.
We first augment the original adjacency matrix $A_{t}$ with self-loops by adding the identity matrix:
$\tilde{A}_{t}=A_{t}+\mathbb{I},$ (2)
where $\mathbb{I}\in\mathbb{R}^{n\times n}$ is the identity matrix. Then, we compute the degree matrix $\tilde{D}_{t}\in\mathbb{R}^{n\times n}$ of $\tilde{A}_{t}$ as:
$\tilde{D}_{t}=\operatorname{diag}\left(\sum_{j=1}^{n}\tilde{A}_{t}(i,j)\right).$ (3)
The normalized adjacency matrix $\hat{A}_{t}$ is then given by:
$\hat{A}_{t}=\tilde{D}_{t}^{-1/2}\tilde{A}_{t}\tilde{D}_{t}^{-1/2}.$ (4)
This normalization ensures that neighborhood feature aggregation is appropriately scaled and robust to variations in node degree. The resulting matrix $\hat{A}_{t}$ acts as a localized spectral filter, approximating graph convolution via eigenspace regularization in a computationally efficient manner.
At time $t$, the input to the first layer is a randomly initialized node feature matrix $F_{t}^{0}\in\mathbb{R}^{n\times d}$, where each row $f_{t}^{i}\in\mathbb{R}^{d}$ represents the latent embedding of node $v_{i}$. This initialization is adopted due to the lack of observed node attributes in the dynamic network.
The output of the final GCN layer at time $t$, denoted as $F_{t}^{l}$, encodes both local topology and structural semantics within the snapshot $g_{t}$. Across time, the set of snapshot-specific node embeddings is represented as:
$F_{t}=\left\{f_{t}^{1},f_{t}^{2},\ldots,f_{t}^{n}\right\},\quad 1\leq t\leq T.$ (5)
3.3.2 Weight Parameter Sharing
Given the dynamic nature of the network, dynamic update strategies are incorporated into the graph convolutional layers. This is necessary because any change in nodes or edges triggers an adjustment of the weight parameters in the convolutional layers to adapt to the evolving network structure. Therefore, we follow the method of EvolveGCNs[15] and use GRU to update the weight parameters of GCNs at the previous time step, which serves as the weight of GCNs at the current time step, to adapt to the development of dynamic networks. The weight update strategy for the $l$-th layer at time t is as follows:
$Z_{t}^{l}=\operatorname{sigmoid}\left(U_{Z}^{l}W_{t-1}^{l}+V_{Z}^{l}F_{t}^{l}+B_{Z}^{l}\right),$ (6)
$R_{t}^{l}=\operatorname{sigmoid}\left(U_{R}^{l}W_{t-1}^{l}+V_{R}^{l}F_{t}^{l}+B_{R}^{l}\right),$
$\hat{W}_{t}^{l}=\tanh\left(U_{W}^{l}\left(R_{t}^{l}\circ W_{t-1}^{l}\right)+V_{W}^{l}F_{t}^{l}+B_{W}^{l}\right),$
$W_{t}^{l}=\Phi\left(F_{t}^{l},W_{t-1}^{l}\right)=\left(1-Z_{t}^{l}\right)\circ\hat{W}_{t}^{l}+Z_{t}^{l}\circ W_{t-1}^{l},$
where $Z_{t}^{l}$ is the the update gate output matrix, $R_{t}^{l}$ the reset gate output matrix and $\hat{W}_{t}^{l}$ the pre-output matrix. $U$ and $V$ are the weight matrix in GRU, and $B$ is a bias term.
The GCNs module aggregates information between nodes and their neighbors, while the GRU adaptively adjusts weight parameters over time. Together, these modules dynamically capture the evolving neighborhood information of the nodes. Thus, operations in the spatial feature extraction stage can be formally represented as
$F_{t}^{l}=\Gamma\left(F_{t}^{l-1},A_{t},W_{t}^{l-1}\right)=\Gamma\left(F_{t}^{l-1},A_{t},\Phi\left(F_{t}^{l-1},W_{t}^{l-1}\right)\right),$ (7)
where $\Gamma(\cdot)$ represents graph convolution operation, and $\Phi(\cdot)$ represents weight update operation using GRU. The spatial feature extraction model is shown in Fig. 2.
Figure 2: Spatial feature extraction stage of GTENN, where GCNs learn node features by aggregating structural information.
3.4 Time Feature Extraction
Classic statistical models and machine learning models can effectively identify sequence data but cannot solve the problem of long-range dependencies of dynamic nodes. That is, in the process of node evolution, the connections established between nodes in the current-time network may have been established a long time ago and only re-established in the current time. In order to obtain information on the dynamic changes of nodes in the time domain and solve the problem of long-distance dependencies between dynamic nodes, we use GRU’s gating mechanism to adjust the input and memory information of temporal data. Compared with LSTM, GRU, instead of using unit states, uses hidden states to transmit information. Therefore, by incorporating GRU in the learning process, the model becomes easier to train, significantly enhancing learning efficiency while achieving comparable results. The structure diagram of GRU is shown in Fig. 3. In the GRU computation process, there are two gating mechanisms, the update gate and the reset gate. The update gate controls the extent to which state information from the previous time step is incorporated into the current state. A larger value indicates that more information on the previous state is retained. Conversely, the reset gate controls the degree to which the previous state is ignored; a smaller value means more emphasis on ignoring the previous state information.
The current temporal feature $F_{t}$ and the hidden state $h_{t-1}$ transmitted from the previous time node form reset and update gates. Among them, the reset gate mainly determines the selection of node information for long-term memory, and it can be obtained as
$r_{t}=\sigma\left(W_{r}F_{t}+U_{r}h_{t-1}+B_{r}\right).$ (8)
The update gate mainly updates the status
$z_{t}=\sigma\left(W_{z}F_{t}+U_{z}h_{t-1}+B_{z}\right).$ (9)
After obtaining the two gate control signals, the reset gate is first applied to select the relevant information from the previous hidden state and the current input data. The information from the two stats is then combined to compute the new memory state
$\hat{h}_{t}=\tanh\left(r_{t}\circ U_{h}h_{t-1}+W_{h}X_{t}+B_{h}\right).$ (10)
where $W$ and $U$ are weight vectors, and $B$ is the bias term. Finally, the update gate selectively retains information from the current hidden state while forgetting the previous hidden state, yielding the current-time feature. The calculation process is as follows
$h_{t}=\left(1-z_{t}\right)\circ h_{t-1}+z_{t}\circ\hat{h}_{t}.$ (11)
Figure 3: Structure of the GRU used for temporal feature extraction in GTENN.
3.5 Community Detection
The algorithm proposed in this paper learns node representations of dynamic networks in an unsupervised manner. To ensure that embeddings reflect the underlying graph topology, we adopt a distance-preserving objective based on pairwise proximity: connected nodes should be embedded close to each other, while unconnected nodes should be placed farther apart. This promotes both first-order (direct link) and second-order (neighborhood structure) proximity preservation.
Let $\mathcal{S}=\{(f_{x},f_{y})\mid(v_{x},v_{y})\in E\}$ be the set of embedding pairs for connected nodes. The pairwise ranking-based loss function is defined as:
$L=\sum_{(f_{x},f_{y})\in\mathcal{S}}\sum_{(f_{x},f_{u})\in\mathcal{S}}\left[m+\|f_{x}-f_{y}\|_{2}^{2}-\|f_{x}-f_{u}\|_{2}^{2}\right]_{+},$ (12)
where $\cdot]_{+}=\max(\cdot,0)$ denotes the hinge loss function, and $m>0$ is a predefined margin constant that reflects the security domain—that is, the minimum separation distance we wish to enforce between a positive (connected) and negative (unconnected) sample. A larger value of $m$ encourages embeddings of unrelated nodes to be pushed further apart.
In large-scale graphs, evaluating all unconnected node pairs is computationally prohibitive. To alleviate this, we adopt the negative sampling[15, 17] strategy. Instead of exhaustively computing losses over all non-edges, we sample a small number of unconnected node pairs as negative examples for each positive pair $(v_{x},v_{y})$. Specifically, negative samples $v_{q}$ are drawn from a noise distribution $P_{n}(v)$, which we define proportionally to node degree: $P_{n}(v)\sim d_{v}^{3/4}$. The modified loss function becomes:
$L=\sum_{(f_{x},f_{y})\in\mathcal{S}}\left(\sum_{q=1}^{Q}\mathbb{E}_{v_{q}\sim P_{n}(v)}\left[m+\|f_{x}-f_{y}\|_{2}^{2}-\|f_{x}-f_{v_{q}}\|_{2}^{2}\right]_{+}\right),$ (13)
where $Q$ is the number of negative samples drawn per positive pair.
After learning the spatial-temporal embeddings $f_{t}^{i}$ of nodes in each snapshot, we apply the Self-Organizing Map (SOM) to discover latent community structure. SOM is an unsupervised neural clustering algorithm that maps high-dimensional vectors into a low-dimensional (typically 2D) grid while preserving topological relations. Using competitive learning, SOM identifies clusters of nodes with similar structural embeddings, forming the final community partitions. The full procedure is summarized in Algorithm 1.
Algorithm 1 Dynamic Community Discovery Method Based on Spatiotemporal Graph Embedding
$G=\{g_{1},g_{2},\dots,g_{t},\dots,g_{T}\}$ $\triangleright$ Input graph sequence
$C=\{C_{1},C_{2},\dots,C_{T}\}$ $\triangleright$ Output community sequence
for $g_{t}$ in $G$ do
  Create time snapshot adjacency matrix $A_{t}\in\mathbb{R}^{n\times n}$
  Create time snapshot unit matrix $\mathbb{I}_{t}\in\mathbb{R}^{n\times n}$
  Initialize node feature matrix $F_{t}\in\mathbb{R}^{n\times d}$
  $W_{t}^{1}\leftarrow\Phi(F_{t},W_{t-1}^{1})$
  $F_{t}^{1}\leftarrow\Gamma(A_{t},F_{t},W_{t}^{1})$
  $W_{t}^{2}\leftarrow\Phi(F_{t}^{1},W_{t-1}^{2})$
  $F_{t}^{2}\leftarrow\Gamma(A_{t},F_{t}^{1},W_{t}^{2})$
  $F_{s}\leftarrow F_{t}^{2}$
  $F_{t}\leftarrow\text{GRU}(F_{s},F_{t-1})$
  $C_{t}\leftarrow\text{SOM}(F_{t})$
  Append $C_{t}$ to $C$
end forreturn $W_{t}^{1},W_{t}^{2},F_{t}$
4 Results and Analysis
4.1 Experiment Setup
All algorithms in this study are implemented in Python 3.10, with the GCN and GRU components developed using PyTorch 2.4.1. Experiments are conducted on a Hygon C86 32-core CPU machine with a Kylin 10 Linux-based operating system, equipped with 512 GB of memory and 8 Hygon DCU GPUs.
The datasets used in this study consist of time-snapshot dynamic networks, including both synthetic and real-world network datasets. For the synthetic data, we adopt the widely used LFR (Lancichinetti–Fortunato–Radicchi) benchmark model[19], which generates graphs with built-in community structure. In this simulation: $N=1000$ specifies the total number of nodes in the graph; $s=9$ defines the number of time snapshots (i.e., network evolution stages); and $\mu\in[0.1,0.8]$ is the mixing parameter, which controls the community strength: lower values of $\mu$ indicate clearer communities (most edges remain within communities), while higher values increase the inter-community connections, making communities harder to detect. By varying $\mu$, we generate eight synthetic dynamic network datasets labeled from LFR1 to LFR8. These provide different levels of community detectability over time.
To evaluate our GTENN model, we also apply it to four real-world dynamic networks of varying size and complexity: PS network representing face-to-face interactions in educational settings [20], CSMN network capturing online social interactions in college environments [21], DBLP collaboration network from computer science publications [20], and Brain networks derived from neuroimaging data [22]. Cluster indicators were used as evaluation criteria in the experiment, mainly including purity, standardized mutual information, heterogeneity, and homogeneity.
The main parameters that need to be estimated in this algorithm include the number of GCN layers Land the weight matrices $W^{l}_{t}$ for each layer, the GRU weight matrices $U$ and $V$ along with bias matrices $B$, the embedding dimension $d$, the number of competing units in the SOM network along with neighborhood range and learning rate, and the number of negative samples $Q$ and security domain size $m$. These parameters are learned through the training process, where the neural network parameters are optimized via backpropagation while the SOM parameters are configured based on competitive learning principles.
4.2 Evaluation Metrics
The experimental evaluation employs four clustering quality metrics to assess community discovery performance:
Purity: Measures the proportion of correctly classified nodes within each cluster, calculated as
$\text{Purity}=\frac{1}{N}\sum_{k=1}^{K}\max_{j}|C_{k}\cap T_{j}|$ (14)
where $C_{k}$ represents the $k$-th discovered community and $T_{j}$ represents the $j$-th true community. Higher purity values indicate better clustering quality.
Normalized Mutual Information (NMI): Quantifies the agreement between discovered and ground-truth communities, defined as
$\text{NMI}=\frac{2\cdot I(C,T)}{H(C)+H(T)}$ (15)
where $I(C,T)$ is the mutual information and $H(\cdot)$ denotes entropy. NMI ranges from 0 to 1, with values closer to 1 indicating better performance.
Homogeneity: Measures whether each cluster contains only members of a single class, ensuring internal consistency within discovered communities.
Completeness: Evaluates whether all members of a given class are assigned to the same cluster, assessing the ability to group related nodes together.
4.3 Datasets
In this experiment, synthetic and real networks are used to comprehensively evaluate algorithm performance. The LFR benchmark networks provide controllable community structures for systematic testing. Real datasets include: PS, CSMN, Brain and DBLP. The summary of these datasets structures can be found in 2
4.3.1 Synthetic Datasets
We used the LFR benchmark to generate dynamic artificial datasets. The LFR benchmark is widely recognized as a standard for evaluating community detection algorithms because it can generate networks with controllable community structure and realistic properties. The following parameters were used:
  • $N=1000$: Total number of nodes in the network
  • $s=9$: Number of time snapshots in the dynamic sequence
  • $\mu\in[0.1,0.8]$: Mixing parameter that controls network complexity
  • Average degree: $\langle k\rangle=15$
  • Maximum degree: $k_{max}=30$
  • Community size range: $[10,50]$ nodes per community
  • Power-law exponent for degree distribution: $\gamma=2.5$
  • Power-law exponent for community size distribution: $\beta=1.5$
The mixing parameter $\mu$ determines the fraction of edges that connect nodes to communities other than their own, where smaller $\mu$ values indicate clearer community structure and larger $\mu$ values represent more ambiguous community boundaries. By varying only the value of $\mu$, we obtained 8 dynamic artificial datasets denoted as LFR1 ($\mu=0.1$), LFR2 ($\mu=0.2$), LFR3 ($\mu=0.3$), LFR4 ($\mu=0.4$), LFR5 ($\mu=0.5$), LFR6 ($\mu=0.6$), LFR7 ($\mu=0.7$), and LFR8 ($\mu=0.8$), corresponding to different levels of community structure clarity from very clear to highly ambiguous.
4.3.2 Real-world Datasets
To validate the performance of GTENN on real-world scenarios, we employed four types of dynamic networks with varying scales and characteristics:
PS (Primary School) Dataset: This dataset records the face-to-face interactions between students and teachers in a primary school in Lyon, France, collected over two days using wearable sensors. The network contains 242 nodes (students and teachers) and evolves over 125 time steps with 20-second intervals. The dataset captures the natural formation and dissolution of social groups in an educational environment, making it ideal for studying temporal community dynamics.
CSMN (College Social Message Network) Dataset: This dataset represents online social interactions among college students over a semester period. It contains 1,899 nodes representing students and 59,835 temporal edges representing various types of interactions including messages, wall posts, and friendship formations. The network evolves over 193 time steps, each representing one day. This dataset is characterized by periodic patterns corresponding to academic schedules and social events.
DBLP Dataset: This is a collaboration network derived from the DBLP computer science bibliography database, focusing on publications from 2000 to 2005. The network contains 12,334 authors as nodes and co-authorship relationships as edges. The temporal evolution spans 6 years, with each time step representing one year. The dataset exhibits clear research community structures that evolve as researchers change their collaboration patterns and research interests.
Brain Dataset: This dataset represents functional brain networks derived from fMRI data, where nodes correspond to brain regions and edges represent functional connectivity between regions. The network contains 264 nodes (brain regions) and evolves over 200 time steps, each representing a different time window during a cognitive task. This dataset is particularly challenging due to its high-dimensional nature and complex temporal dynamics.
Dataset Nodes Time Steps Avg. Edges/Step Domain
LFR1-8 1,000 9 $\sim$7,500 Synthetic
PS 242 125 $\sim$400 Social
CSMN 1,899 193 $\sim$310 Social
DBLP 12,334 6 $\sim$25,000 Collaboration
Brain 264 200 $\sim$1,200 Biological
Table 2: Statistics of the dynamic network datasets used in our experiments. Each dataset is characterized by the number of nodes, time steps, average edges per step, and its corresponding domain.
4.4 Comparison Algorithms
In order to evaluate the ability of the algorithm for community discovery in this article, it is mainly compared with dynamic community discovery algorithms Dyperm and FacetNet, static community discovery algorithm Louvain, dynamic community discovery method DyCTWE based on time walk graph embedding, and GTENN-K using K-means as the final clustering method. This algorithm is referred to as GTENN in this paper, and purity and NMI are used to evaluate the quality of clustering results.
The evaluation of spatiotemporal graph embedding models is performed by comparing them with traditional static graph embedding methods SDNE[23], DeepWalk[24] and dynamic graph embedding algorithms DynAE, DynGEM[25]. After obtaining node embeddings, K-means clustering is used to divide the nodes into communities. In this paper, purity and normalized mutual information (NMI) are used to evaluate the quality of the clustering results. The parameter settings for the relevant dynamic graph embedding algorithm are shown in Tab. 3.
Algorithm Parameter settings
DynAE $\beta=5,lb=2$
DynGEM $\beta=5,a=1e{-}5,v_{1}=v_{2}=1e{-}6,$
$\rho=0.3,\theta=0.01$
DeepWalk $l=40,w=10,\gamma=6$
Table 3: Parameter settings of related algorithms.
To evaluate the community discovery capability of the proposed algorithm, we compared the GTENN-K, where K stands for the final K-means clustering, with dynamic community discovery algorithms Dyperm and FacetNet[26], the static community discovery algorithm Louvain[27], and the dynamic community discovery method DyCTWE based on time walk graph embedding.
4.5 Results
4.5.1 Ablation Experiments
In complex machine learning models, which often consist of multiple modules, ablation experiments are used to assess the importance of specific components by removing certain modules and rerunning the experiments. The results are then compared with those from the original model to evaluate the impact of the removed modules. We evaluate the effectiveness of dynamic community methods based on spatiotemporal graph embedding through ablation experiment. This experiment is conducted on a dynamic artificial network (LFR2) with $\mu=0.2$. We compare the full model with GRU, GCNs, and GCNs-GRU. The evaluation criteria include the average NMI, completeness, and homogeneity of the dynamic network, which are used to verify the superiority of the proposed algorithm. A detailed comparison is presented in Tab 4.
As shown in Tab. 4, the proposed GTENN consistently outperforms its individual components, demonstrating its effectiveness in capturing the complex spatiotemporal relationships of dynamic networks. GTENN achieves superior performance across all evaluation metrics, including NMI, completeness, and homogeneity, confirming that its integrated design is well-suited for dynamic community discovery. The relatively low performance of the GRU baseline can be attributed to its emphasis on temporal dynamics while neglecting spatial structural characteristics. Without explicit modeling of network topology, GRU struggles to capture node dependencies, leading to weaker community detection. In addition, the random initialization of node features without fixed semantics introduces inconsistencies in representation, further diminishing its effectiveness. Furthermore, the GCNs-GRU hybrid achieves substantially better results than GCNs alone, highlighting the importance of jointly modeling spatial topology and temporal evolution. This synergy between GCNs and GRU provides a more comprehensive representation of evolving communities and yields consistently higher performance across all metrics.
Algorithm NMI Comp Homo
GCNs 0.748226 0.748214 0.748249
GRU 0.705708 705697 0.705726
GCNs-GRU 0.902130 0.900143 0.902308
GTENN 0.948902 0.948017 0.949921
Table 4: Comparison of ablation experimental results on LFR2 dataset ($\mu=0.2$). The table demonstrates the effectiveness of different component combinations in GTENN
4.5.2 Comparative Analysis of Embedding Methods
In Fig. 4 we present the evaluation results of various graph embedding algorithms on dynamic artificial synthetic datasets and real datasets. We first analyze the performance of the embedding algorithms on synthetic networks.
(a) Average purity on Synthetic Datasets
(b) Average NMI on Synthetic Datasets
(c) Average Purity on Real Datasets
(d) Average NMI on Real Datasets
Figure 4: Performance comparison of graph embedding methods on synthetic and real datasets. Results show that GTENN consistently achieves higher purity and NMI scores compared to static methods (SDNE, DeepWalk) and dynamic methods (DynAE, DynGEM).
From Fig. 4(a) and  4(b), it is observed that as the complexity coefficient increases, the average evaluation results for synthetic networks show a decreasing trend. This indicates that as the complexity of the network grows, the ability of the graph embedding algorithm to represent the network structure declines. The GTENN-K algorithm exhibits strong representation capabilities for artificial networks, with K-means yielding superior community partitioning accuracy measured in purity and NMI compared to other embedding models. This demonstrates that the spatiotemporal graph embedding model effectively captures spatiotemporal relationships and preserves node interaction information in complex dynamic networks. The dynamic graph embedding methods DynGEM and DynAE can also provide good embedding features, and the evaluation results of community partitioning achieve good results when the network complexity is low. However, as complexity increases, the evaluation results rapidly decline. This may be attributed to the fact that these two algorithms primarily focus on learning the spatial structure of nodes during graph embedding. For instance, DynGEM only considers the influence of the previous time snapshot on the current one, neglecting the long-range dependencies of dynamic nodes. Consequently, as network complexity increases, the graph embedding performance declines, impacting community discovery effectiveness. The static embedding algorithm DeepWalk achieves strong results, while SDNE shows stable but modest performance on artificial datasets, with minimal changes in NMI values. This suggests that SDNE can consistently characterize dynamic artificial networks, but its ability to extract node features requires further improvement.
We then analyze the performance of the embedded model in real networks and present the results in Fig. 4(c) and  4(d).
From Fig. 4(c) and  4(d), it is evident that all graph embedding models perform better in node partitioning accuracy measured in purity and NMI on smaller dynamic datasets than on larger ones. The decrease in purity and NMI as the network size grows can be attributed to the increase in the number of nodes and the more complex interaction relationships between them. Larger networks introduce additional challenges, such as denser connections, overlapping communities, and a greater diversity of dynamic interactions, which make it harder for models to effectively capture and preserve all relevant spatiotemporal features.
Despite these challenges, the GTENN proposed in this paper consistently maintains high accuracy measured in purity and NMI across both small and large datasets. This highlights its robustness and its ability to effectively extract dynamic spatiotemporal information, allowing it to characterize the evolution of real-world networks more accurately. The model’s ability to integrate spatial and temporal embeddings enables it to uncover subtle patterns and interactions that are often missed by other methods.
Among the baseline models, DeepWalk achieves the second-best performance due to its ability to learn meaningful node representations through random walks. However, its lack of temporal modeling limits its performance on dynamic datasets. SDNE, on the other hand, yields relatively lower purity and NMI on large-scale dynamic datasets, probably because its emphasis on preserving local structures struggles to scale effectively with the increasing complexity of larger networks. Among dynamic graph embedding algorithms, DynGEM outperforms DynAE in community discovery, highlighting that SDNE’s embedding performance on the network surpasses that of fully connected neural networks.
4.5.3 Comparative Analysis of Community Discovery Methods
In Fig. 5(a) and  5(d) we present the evaluation results of community discovery algorithms on dynamic networks.
(a) Average purity on Synthetic Datasets
(b) Average NMI on Synthetic Datasets
(c) Average Purity on Real Datasets
(d) Average NMI on Real Datasets
Figure 5: Comprehensive comparison of community discovery algorithms. GTENN demonstrates superior performance over traditional static (Louvain), dynamic (Dyperm, FacetNet), and embedding-based (DyCTWE) methods across both synthetic and real datasets.
From these figures, it is clear that GTENN shows a noticeable effect on community partitioning for both real and synthetic network datasets. However, its performance on large real dynamic networks is significantly lower than on smaller ones, highlighting that the complexity and variability of large dynamic networks remain key challenges in the research. However, GTENN has achieved strong evaluation results across both artificially synthesized and real dynamic network datasets. The evaluation metrics, purity and NMI in various datasets, are presented in Tab. 5.
Dataset Purity (%) NMI (%)
LFR2 93.73 94.89
LFR4 89.84 93.25
LFR6 78.17 86.36
LFR8 24.03 52.31
PS 92.68 90.10
CSMN 86.64 67.45
Brain 46.91 41.56
DBLP 48.11 37.60
Table 5: Purity and NMI in the synthetic and real datasets of the proposed GTENN alogrithm
Comparing the proposed algorithm with traditional static community discovery algorithms, such as Louvain, and dynamic community discovery algorithms such as Dyperm and FacetNet, GTENN outperforms these methods in community partitioning. FacetNet’s performance on artificial networks is suboptimal, although it performs better on real networks, suggesting that it focuses more on changes in real-world social activities. Dyperm demonstrates relatively stable community partitioning performance, while Louvain performs well in real online communities.
Figure 6: Number of communities identified by GTENN and DyCTWE on the PS dataset, compared with the ground-truth community structure across time steps.
The comparison between GTENN and GTENN-K primarily evaluates the impact of clustering algorithms on the partitioning results of the graph embedding model. From the results, it is observed that the evaluation results for both the SOM and the K-means clustering algorithms are similar, with differences within 5%. The SOM algorithm was chosen because it does not require specifying the number of clusters in advance, only the number of competing units, neighborhood range, and learning rate. This allows for adaptive community division based on the embedding representation of nodes. In contrast, the K-means algorithm requires prior knowledge of the number of clusters, which is often difficult to determine in real dynamic networks. Hence, SOM clustering is more suitable for this work.
(a) Average purity in PS
(b) Average NMI in PS
Figure 7: Evaluation results of GTENN and DyCTWE on the PS dataset over time. The plots report Purity (left) and NMI (right) at each time step, showing that GTENN maintains consistently higher performance across the dynamic snapshots.
In Fig. 6, we present the comparison of community partitioning results for the PS dataset, while in Fig. 7, we present the evaluation results of both algorithms on the same dataset. From Fig. 6, it is evident that GTENN’s partitioning results are closely aligned with the actual number of communities, with only minor discrepancies. In contrast, DyCTWE shows a maximum difference of 5 communities from the real partitioning. Fig. 7 reveals that the quality of DyCTWE’s partitioning deteriorates over time, indicating error accumulation and a decline in community discovery performance. This issue likely arises because DyCTWE overly emphasizes local incremental updates, neglecting the overall evolution of the network. On the other hand, GTENN demonstrates an increasing trend in performance over time, suggesting that spatiotemporal graph embedding effectively captures the dynamic evolution of nodes and offers robust representation capabilities for dynamic networks.
4.6 Comparison with Attention-Based Methods
In this section, we present the evaluation results in terms of purity and NMI and computing power consumption agains some attention-based algorithms, including ASTGCN [28], MSTGCN [29] and ST-GAT [30].
The results in Table 6 demonstrate that the proposed GTENN algorithm exhibits outstanding performance in both community detection quality and computational efficiency across two representative datasets: LFR2 and PS.
Table 6: Comprehensive Performance Comparison of Different Algorithms
Algorithm Dataset Purity NMI Training Time Memory/Params
GTENN LFR2 0.937 0.949 12.3 2.1GB/48K
PS 0.927 0.901 8.5 1.8GB/48K
NWFCM LFR2 0.721 0.698 6.2 0.8GB/12K
PS 0.704 0.681 4.1 0.6GB/12K
ASTGCN LFR2 0.834 0.821 28.6 4.5GB/156K
PS 0.798 0.785 18.7 3.8GB/156K
MSTGCN LFR2 0.856 0.843 45.3 6.8GB/243K
PS 0.841 0.827 32.1 5.9GB/243K
ST-GAT LFR2 0.812 0.795 22.4 3.7GB/128K
PS 0.789 0.773 16.2 3.1GB/128K
The proposed GTENN achieves Purity values of 0.937 and 0.927 on the LFR2 and PS datasets, respectively, with corresponding NMI values of 0.949 and 0.901, significantly outperforming all baseline algorithms. Compared to the second-best performing MSTGCN, GTENN shows an average improvement of approximately 8.2% in Purity and 9.1% in NMI. This indicates that GTENN, through its spatiotemporal joint modeling strategy combining GCN and GRU, can more accurately capture community evolution patterns in dynamic networks than attention-based approaches.
The performance comparison reveals distinct characteristics of different algorithm categories:
  • Traditional methods like NWFCM [31] show the lowest performance but maintain computational efficiency.
  • Attention-based methods (ASTGCN, MSTGCN, ST-GAT) achieve moderate performance but with significant computational overhead.
  • Our spatiotemporal approach achieves the best performance while maintaining reasonable computational cost.
GTENN also demonstrates clear efficiency advantages over attention-based methods. Although its training time is slightly higher than that of the traditional NWFCM approach, it remains substantially lower than that of more complex attention-based models. Specifically, compared with MSTGCN, GTENN reduces training time by approximately 70
The robustness of GTENN is further demonstrated by its stable performance across datasets of different nature and complexity. From the synthetic LFR2 to the real-world PS network, the decline in Purity is less than 2%, while attention-based algorithms typically experience 3-8% decreases. This demonstrates the superior generalization capability and stability of the proposed algorithm. The consistent performance across different dataset characteristics (synthetic vs. real-world, different scales and domains) indicates that our spatiotemporal modeling approach is more robust than attention mechanisms that may be sensitive to specific network structures.
The results further underscore the advantages of GTENN over attention-based methods. Specifically, ASTGCN, although effective for traffic prediction, shows limited suitability for community detection tasks. MSTGCN leverages multi-view attention mechanisms but incurs substantial computational overhead, making it less practical for large or complex networks. ST-GAT integrates graph attention networks for spatiotemporal modeling, yet its performance exhibits instability across different datasets. In conclusion, the proposed GTENN achieves superior detection accuracy measured in purity and NMI, and computational efficiency, offering more stable and effective spatiotemporal modeling than attention-based counterparts, while maintaining substantially lower computational complexity.
5 Conclusion
In this work, we proposed GTENN, a community discovery algorithm based on spatiotemporal graph embedding, to explore the complex structure and dynamics of social networks. Experimental validation demonstrates that GTENN effectively uncovers community evolution in dynamic networks, providing new insights into group organization and information flow.
A key characteristic of GTENN is its discrete-time snapshot framework, which is well suited to datasets collected in aggregated intervals. While continuous-time approaches such as TGN and JODIE offer fine-grained event-level modeling, our datasets are not inherently event-driven, and converting them into event streams could distort their temporal structure. For this reason, we adopt the snapshot-based representation, while extending GTENN to continuous-time settings with appropriate datasets remains an important direction for future research.
Looking ahead, we plan to enhance GTENN with advanced deep learning architectures, such as attention-based networks, to strengthen spatiotemporal feature learning and improve performance across evaluation metrics. We will also extend the framework to diverse domains, including transportation, biology, and particle interactions, to test its generality. Of particular significance is the financial sector, where applications in fraud detection and anti–money laundering (AML) can uncover hidden transactional patterns and anomalous group behaviors that traditional methods often miss. Such extensions would not only advance domain-specific impact but also demonstrate the broader practical utility of GTENN for understanding complex and evolving networks.
Acknowledgment
The authors gratefully acknowledge the support of Qilu Bank, which provided the necessary CPU and GPU machine time for the computational experiments. We also wish to express our gratitude to its Financial Mathematical Lab for providing a conducive research environment and valuable resources.
Data Availability Statement
This study utilized publicly available open datasets, including real-world dynamic network data and benchmark datasets. The synthetic datasets generated and analyzed during the study are available from the corresponding author upon reasonable request.